the sum of A and B is 30 years and 5 years hence A will be three times as old as B. Find their ages.
Answer by stanbon(75887)

the sum of A and B is 30 years and 5 years hence A will be three times as old as B. Find their ages.
--------------------
A + B = 30
A+5 = 3(B+5)
---------------------
Substitute for "A" and solve for "B"::
30-B + 5 = 3B+15
------
4B = 20
B = 5 yrs
A = 25 yrs
-----------------
Cheers,
Stan H.
------------